# MM-LLMs: Recent Advances in MultiModal Large Language Models

### 25 Jan 2024 v2

@@ 3.1 MM PT
During the PT stage, typically leveraging the X-Text datasets, Input and Output Projectors are trained to achieve alignment among various modalities by optimizing predefined objectives (PEFT is sometimes applied to the LLM Backbone). For MM understanding models, optimization focuses solely on Equation (2), while for MM generation models, optimization involves Equations (2), (4), and (5). In the latter case, Equation (2) also includes the ground-truth signal token sequence.
@@
The X-Text datasets encompass Image-Text, Video-Text, and Audio-Text, with Image-Text having two types: Image-Text pairs (i.e., <img1><txt1>) and interleaved Image-Text corpus (i.e., <txt1><img1><txt2><txt3><img2><txt4>). The detailed statistics for these X-Text datasets are presented in Table 3 of Appendix F.
@@ 3.2 MM IT
MM IT is a methodology that entails the fine-tuning of pre-trained MM-LLMs using a set of instruction-formatted datasets Wei et al. (2021). Through this tuning process, MM-LLMs can generalize to unseen tasks by adhering to new instructions, thereby enhancing zero-shot performance. This straightforward yet impactful concept has catalyzed the success of subsequent endeavors in the field of NLP, exemplified by works such as InstructGPT Ouyang et al. (2022), OPT-IML Iyer et al. (2022), and InstructBLIP Dai et al. (2023).
@@
MM IT comprises Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), aiming to align with human intents or preferences and enhance the interaction capabilities of MM-LLMs. SFT converts part of the PT stage data into an instruction-aware format. Using visual Question-Answer (QA) as an example, various templates may be employed like (1) <Image>{Question} A short answer to the question is; (2) <Image>Examine the image and respond to the following question with a brief answer: {Question}. Answer:; and so on. Next, it fine-tunes the pre-trained MM-LLMs using the same optimization objectives. The SFT dataset can be structured as either single-turn QA or multi-turn dialogues.
@@
After SFT, RLHF involves further fine-tuning of the model, relying on feedback regarding the MM-LLMs’ responses (e.g., Natural Language Feedback (NLF) labeled manually or automatically) Sun et al. (2023). This process employs a reinforcement learning algorithm to effectively integrate the non-differentiable NLF. The model is trained to generate corresponding responses conditioned on the NLF Chen et al. (2023h); Akyürek et al. (2023). The statistics for SFT and RLHF datasets are presented in Table 4 of Appendix F.
The datasets used by existing MM-LLMs in the MM PT and MM IT stages are diverse, but they are all subsets of the datasets in Tables 3 and 4.
| Flamingo | I+V+TT | I/V: NFNet-F6 | Cross-attention | Chinchilla-1.4B/7B/70B (Frozen) | – | – | – | – |
| BLIP-2 | I+TT | I: CLIP/Eva-CLIP ViT@224 | Q-Former w/ Linear Projector | Flan-T5/OPT (Frozen) | – | – | 129M | – |
| LLaVA | I+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-7B/13B (PT: Frozen; IT: PEFT) | – | – | – | – |
| MiniGPT-4 | I+TT | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-13B (PT: Frozen; IT: PEFT) | – | – | – | – |
| mPLUG-Owl | I+TT | I: CLIP ViT-L/14 | Cross-attention | LLaMA-7B(PT: Frozen; IT: PEFT) | – | – | – | – |
| X-LLM | I+V+A+TT | I/V: ViT-G; A: C-Former | Q-Former w/ Linear Projector | ChatGLM-6B (Frozen) | – | – | – | – |
| VideoChat | V+TT | I: ViT-G | Q-Former w/ Linear Projector | Vicuna (Frozen) | – | – | – | – |
| InstructBLIP | I+V+TT | I/V: ViT-G/14@224 | Q-Former w/ Linear Projector | Flan-T5/Vicuna (Frozen) | – | – | 129M | 1.2M |
| PandaGPT | I+TT | I: ImageBind | Linear Projector | Vicuna-13B (PEFT) | – | – | – | – |
| PaLI-X | I+TT | I: ViT | Linear Projector | UL2-32B (PEFT) | – | – | – | – |
| Video-LLaMA | I+V+A+TT | I/V: EVA-CLIP ViT-G/14; A: ImageBind | Q-Former w/ Linear Projector | Vicuna/LLaMA (Frozen) | – | – | – | – |
| Video-ChatGPT | V+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-v1.1 (Initialized with LLaVA, Frozen) | – | – | – | – |
| Shikra | I+TT | I: CLIP ViT-L/14@224 | Linear Projector | Vicuna-7B/13B (PEFT) | – | – | 600K | 5.5M |
| DLP | I+TT | I: CLIP/Eva-CLIP ViT | Q-Former+P-Former w/ Linear Projector | OPT/Flan-T5 (Frozen) | – | – | – | – |
| BuboGPT | I+A+TT | I: CLIP/Eva-CLIP ViT; A: ImageBind | Q-Former w/ Linear Projector | Vicuna (Frozen) | – | – | – | – |
| ChatSpot | I+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-7B/LLaMA (PT: Frozen; IT: PEFT) | – | – | – | – |
| Qwen-VL-(Chat) | I+TT | I: ViT@448 initialized from OpenClip’s ViT-bigG | Cross-attention | Qwen-7B (PT: Frozen; IT: PEFT) | – | – | 1.4B | 50M |
| NExT-GPT | I+V+A+TI+V+A+T | I/V/A: ImageBind | Linear Projector | Vicuna-7B (PEFT) | Tiny Transformer | I: Stable Diffusion; V: Zeroscope; A: AudioLDM | – | – |
| MiniGPT-5 | I+TI+T | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-7B (PEFT) | Tiny Transformer w/ MLP | I: StableDiffusion-2 | – | – |
| LLaVA-1.5 | I+TT | I: CLIP ViT-L@336 | MLP | Vicuna-v1.5-7B/13B (PT: Frozen; IT: PEFT) | – | – | 0.6M | 0.7M |
| MiniGPT-v2 | I+TT | I: Eva-CLIP ViT@448 | Linear Projector | LLaMA-2-Chat-7B (PEFT) | – | – | – | – |
| CogVLM | I+TT | I: Eva-2-CLIP ViT | MLP | Vicuna-v1.5-7B (PEFT) | – | – | – | – |
| DRESS | I+TT | I:Eva-CLIP ViT-G/14 | Linear Projector | Vicuna-v1.5-13B (PEFT) | – | – | – | – |
| X-InstructBLIP | I+V+A+3D+TT | I/V: Eva-CLIP ViT-G/14; A: BEATs; 3D: ULIP-2 | Q-Former w/ Linear Projector | Vicuna-v1.1-7B/13B (Frozen) | – | – | – | – |
| CoDi-2 | I+V+A+TI+V+A+T | I/V/A: ImageBind | MLP | LLaMA-2-Chat-7B (PT: Frozen; IT: PEFT) | MLP | I: Stable Diffusion-2.1; V: Zeroscope-v2; A: AudioLDM-2 | – | – |
| VILA | I+TT | I: ViT@336 | Linear Projector | LLaMA-2-7B/13B (PEFT) | – | – | 50M | 1M |
@@ 4 SOTA MM-LLMs
Based on the previously defined design formulations, we conduct a comprehensive comparison of the architectures and training dataset scales for SOTA MM-LLMs, as illustrated in Table 1. Subsequently, we will provide a concise introduction to the core contributions of these models and summarize their developmental trends.
(1) Flamingo Alayrac et al. (2022) represents a series of Visual Language (VL) Models designed for processing interleaved visual data and text, generating free-form text as the output. (2) BLIP-2 Li et al. (2023c) introduces a more resource-efficient framework, comprising the lightweight Q-Former to bridge modality gaps and the utilization of frozen LLMs. Leveraging LLMs, BLIP-2 can be guided for zero-shot image-to-text generation using natural language prompts. (3) LLaVA Liu et al. (2023e) pioneers the transfer of IT techniques to the MM domain. Addressing data scarcity, LLaVA introduces a novel open-source MM instruction-following dataset created using ChatGPT/GPT-4, alongside the MM instruction-following benchmark, LLaVA-Bench. (4) MiniGPT-4 Zhu et al. (2023a) proposes a streamlined approach where training only one linear layer aligns the pre-trained vision encoder with the LLM. This efficient method enables the replication of the exhibited capabilities of GPT-4. (5) mPLUG-Owl Ye et al. (2023) presents a novel modularized training framework for MM-LLMs, incorporating the visual context. To assess different models’ performance in MM tasks, the framework includes an instructional evaluation dataset called OwlEval. (6) X-LLM Chen et al. (2023b) is expanded to various modalities, including audio, and demonstrates strong scalability. Leveraging the language transferability of the Q-Former, X-LLM is successfully applied in the context of Sino-Tibetan Chinese. (7) VideoChat Li et al. (2023d) pioneers an efficient chat-centric MM-LLM for video understanding dialogue, setting standards for future research in this domain and offering protocols for both academia and industry. (8) InstructBLIP Dai et al. (2023) is trained based on the pre-trained BLIP-2 model, updating only the Q-Former during MM IT. By introducing instruction-aware visual feature extraction and corresponding instructions, the model enables the extraction of flexible and diverse features. (9) PandaGPT Su et al. (2023) is a pioneering general-purpose model with the capability to comprehend and act upon instructions across different modalities: text, image/video, audio, thermal, depth, and inertial measurement units. (10) PaLI-X Chen et al. (2023g) is trained using mixed VL objectives and unimodal objectives, including prefix completion and masked-token completion. This approach proves effective for both downstream task results and achieving the Pareto frontier in the fine-tuning setting. (11) Video-LLaMA Zhang et al. (2023e) introduces a multi-branch cross-modal PT framework, enabling LLMs to simultaneously process the vision and audio content of a given video while engaging in conversations with humans. This framework aligns vision with language as well as audio with language. (12) Video-ChatGPT Maaz et al. (2023) is a model specifically designed for video conversations, capable of generating discussions about videos by integrating spatiotemporal vision representations. (13) Shikra Chen et al. (2023d) introduces a simple and unified pre-trained MM-LLM tailored for Referential Dialogue, a task involving discussions about regions and objects in images. This model demonstrates commendable generalization ability, effectively handling unseen settings. (14) DLP Jian et al. (2023) proposes the P-Former to predict the ideal prompt, trained on a dataset of single-modal sentences. This showcases the feasibility of single-modal training to enhance MM learning. (15) BuboGPT Zhao et al. (2023d) is a model constructed by learning a shared semantic space for a comprehensive understanding of MM content. It explores fine-grained relationships among different modalities such as image, text, and audio. (16) ChatSpot Zhao et al. (2023b) introduces a simple yet potent method for finely adjusting precise referring instructions for MM-LLM, facilitating fine-grained interactions. The incorporation of precise referring instructions, consisting of image- and region-level instructions, enhances the integration of multi-grained VL task descriptions. (17) Qwen-VL Bai et al. (2023b) is a multi-lingual MM-LLM that supports both English and Chinese. Qwen-VL also allows the input of multiple images during the training phase, improving its ability to understand the vision context. (18) NExT-GPT Wu et al. (2023d) is an end-to-end, general-purpose any-to-any MM-LLM that supports the free input and output of image, video, audio, and text. It employs a lightweight alignment strategy, utilizing LLM-centric alignment in the encoding phase and instruction-following alignment in the decoding phase. (19) MiniGPT-5 Zheng et al. (2023b) is an MM-LLM integrated with inversion to generative vokens and integration with Stable Diffusion. It excels in performing interleaved VL outputs for MM generation. The inclusion of classifier-free guidance during the training phase enhances the quality of generation.
@@
For introduction regarding the remaining seven MM-LLMs, please refer to Appendix D, which includes (20) LLaVA-1.5 Liu et al. (2023d), (21) MiniGPT-v2 Chen et al. (2023c), (22) CogVLM Wang et al. (2023), (23) DRESS Chen et al. (2023h), (24) X-InstructBLIP Panagopoulou et al. (2023), (25) CoDi-2 Tang et al. (2023a), and (26) VILA Lin et al. (2023).
@@ 5 Benckmarks and Performance
To offer a comprehensive performance comparison, we have compiled a table featuring major MM-LLMs across VL benchmarks gathered from various papers Li et al. (2023c); Chen et al. (2023c, e); Lin et al. (2023), as shown in Table 2. The information of these benchmarks can be found in Appendix E. Next, we will extract essential training recipes that boost the effectiveness of MM-LLMs, drawing insights from SOTA models.
@@
Firstly, higher image resolution can incorporate more visual details for the model, benefiting tasks that require fine-grained details. For example, LLaVA-1.5 and VILA employ a resolution of , while Qwen-VL and MiniGPT-v2 utilize . However, higher resolutions lead to longer token sequences, incurring additional training and inference costs. MiniGPT-v2 addresses this by concatenating adjacent visual tokens in the embedding space to reduce length. Recently, Monkey Li et al. (2023h) proposed a solution to enhance the resolution of input images without retraining a high-resolution visual encoder, utilizing only a low-resolution visual encoder, supporting resolutions up to . To enhance the understanding of rich-text images, tables, and document content, DocPedia Feng et al. (2023) introduced a method to increase the visual encoder resolution to , overcoming the limitations of poorly performing low resolutions in open-sourced ViT. Secondly, the incorporation of high-quality SFT data can significantly improve performance in specific tasks, as evidenced by the addition of ShareGPT4V data to LLaVA-1.5 and VILA-13B, as shown in Table 2. Moreover, VILA reveals several key findings: (1) Performing PEFT on the LLM Backbone promotes deep embedding alignment, crucial for ICL; (2) Interleaved Image-Text data proves beneficial, whereas Image-Text pairs alone are sub-optimal; (3) Re-blending text-only instruction data (e.g., Unnatural instruction Honovich et al. (2022)) with image-text data during SFT not only addresses the degradation of text-only tasks but also enhances VL task accuracy.

### 20 Feb 2024 v4

@@ 3.1 MM PT
During the PT stage, typically leveraging the X-Text datasets, Input and Output Projectors are trained to achieve alignment among various modalities by optimizing predefined objectives. For MM understanding models, optimization focuses solely on Equation (2), while for MM generation models, optimization involves Equations (2), (4), and (5). In the latter case, Equation (2) also includes the ground-truth signal token sequence.
@@
The X-Text datasets include Image-Text, Video-Text, and Audio-Text, with Image-Text having two types: Image-Text pairs (e.g., <img1> <txt1>) and interleaved Image-Text corpus (e.g., <txt1><img1><txt2><txt3><img2><txt4>). Details of X-Text datasets are shown in Table 3 of Appendix G.
| Flamingo | I+V+TT | I/V: NFNet-F6 | Cross-attention | Chinchilla-1.4B/7B/70B | – | – | – | – |
| BLIP-2 | I+TT | I: CLIP/Eva-CLIP ViT@224 | Q-Former w/ Linear Projector | Flan-T5/OPT | – | – | 129M | – |
| LLaVA | I+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-7B/13B | – | – | – | – |
| MiniGPT-4 | I+TT | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-13B | – | – | – | – |
| mPLUG-Owl | I+TT | I: CLIP ViT-L/14 | Cross-attention | LLaMA-7B | – | – | – | – |
| Otter | I+TT | I: CLIP ViT-L/14 | Cross-attention | LLaMA-7B | – | – | – | – |
| X-LLM | I+V+A+TT | I/V: ViT-G; A: C-Former | Q-Former w/ Linear Projector | ChatGLM-6B | – | – | – | – |
| VideoChat | V+TT | I: ViT-G | Q-Former w/ Linear Projector | Vicuna | – | – | – | – |
| InstructBLIP | I+V+TT | I/V: ViT-G/14@224 | Q-Former w/ Linear Projector | Flan-T5/Vicuna | – | – | 129M | 1.2M |
| PandaGPT | I+TT | I: ImageBind | Linear Projector | Vicuna-13B | – | – | – | – |
| GILL | I+TI+T | I: CLIP ViT-L | Linear Projector | OPT-6.7B | Tiny Transformer | I: Stable Diffusion-1.5 | – | – |
| PaLI-X | I+TT | I: ViT | Linear Projector | UL2-32B | – | – | – | – |
| Video-LLaMA | I+V+A+TT | I/V: Eva-CLIP ViT-G/14; A: ImageBind | Q-Former w/ Linear Projector | Vicuna/LLaMA | – | – | – | – |
| Video-ChatGPT | V+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-v1.1 | – | – | – | – |
| Shikra | I+TT+ | I: CLIP ViT-L/14@224 | Linear Projector | Vicuna-7B/13B | – | – | 600K | 5.5M |
| LLaVAR | I+TT | I: CLIP ViT-L/14@224 & CLIP ViT-L/14@336 | Linear Projector | Vicuna-13B | – | – | – | – |
| mPLUG-DocOwl | +TT | I: CLIP ViT-L/14 | Cross-attention | LLaMA-7B | – | – | – | – |
| Lynx | I+V+TT | I/V: Eva-CLIP ViT-1B | Cross-attention | Vicuna | – | – | – | – |
| Emu | I+V+TI+T | I/V: Eva-CLIP-1B | Cross-attention | LLaMA-13B | MLP | I: Stable Diffusion-1.5 | – | – |
| DLP | I+TT | I: CLIP/Eva-CLIP ViT | Q-Former+P-Former w/ Linear Projector | OPT/Flan-T5 | – | – | – | – |
| BuboGPT | I+A+TT+ | I: CLIP/Eva-CLIP ViT; A: ImageBind | Q-Former w/ Linear Projector | Vicuna | – | – | – | – |
| ChatSpot | I+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-7B/LLaMA | – | – | – | – |
| IDEFICS | I+TT | I: OpenCLIP | Cross-attention | LLaMA | – | – | – | – |
| Qwen-VL-(Chat) | I+TT | I: ViT@448 initialized from OpenClip’s ViT-bigG | Cross-attention | Qwen-7B | – | – | 1.4B | 50M |
| LaVIT | I+TI+T | I: ViT | Cross-attention | LLaMA-7B | – | I: Stable Diffusion | – | – |
| NExT-GPT | I+V+A+TI+V+A+T | I/V/A: ImageBind | Linear Projector | Vicuna-7B | Tiny Transformer | I: Stable Diffusion; V: Zeroscope; A: AudioLDM | – | – |
| DreamLLM | I+TI+T | I: CLIP ViT-L | Linear Projector | Vicuna | MLP | I: Stable Diffusion | – | – |
| AnyMAL | I+V+A+TT | I: CLIP ViT/L & ViT-G & DinoV2; V: Intervideo; A: CLAP | I/V: Cross-attention; A: Linear Projector | LLaMA-2 | – | – | – | – |
| MiniGPT-5 | I+TI+T | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-7B | Tiny Transformer w/ MLP | I: StableDiffusion-2 | – | – |
| LLaVA-1.5 | I+TT | I: CLIP ViT-L@336 | MLP | Vicuna-v1.5-7B/13B | – | – | 0.6M | 0.7M |
| MiniGPT-v2 | I+TT | I: Eva-CLIP ViT@448 | Linear Projector | LLaMA-2-Chat-7B | – | – | – | – |
| CogVLM | I+TT | I: Eva-2-CLIP ViT | MLP | Vicuna-v1.5-7B | – | – | – | – |
| Qwen-Audio | A+TT | A: Whisper-L-v2 | Linear Projector | Qwen-7B | – | – | – | – |
| DRESS | I+TT | I:Eva-CLIP ViT-G/14 | Linear Projector | Vicuna-v1.5-13B | – | – | – | – |
| X-InstructBLIP | I+V+A+3D+TT | I/V: Eva-CLIP ViT-G/14; A: BEATs; 3D: ULIP-2 | Q-Former w/ Linear Projector | Vicuna-v1.1-7B/13B | – | – | – | – |
| CoDi-2 | I+V+A+TI+V+A+T | I/V/A: ImageBind | MLP | LLaMA-2-Chat-7B | MLP | I: Stable Diffusion-2.1; V: Zeroscope-v2; A: AudioLDM-2 | – | – |
| RLHF-V | I+TT | I: BEiT-3 | Linear Projector | Vicuna-v1-13B | – | – | – | – |
| Silkie | I+TT | I: ViT initialized from OpenCLIP’s ViT-bigG | Cross-attention | Qwen-7B | – | – | – | – |
| Lyrics | I+TT | I: CLIP ViT-L/14 & Grounding-DINO-T w/ Swin-T & SAM-HQ w/ MAE & ViT-H & RAM++ w/ Swin-B | MQ-Former w/ Linear Projection | Vicuna-13B | – | – | – | – |
| VILA | I+TT | I: ViT@336 | Linear Projector | LLaMA-2-7B/13B | – | – | 50M | 1M |
| IntrenVL | I+V+TT | I/V: InternViT-6B; T: LLaMA-7B | Cross-attention w/ MLP | QLLaMA-8B & Vicuna-13B | – | – | – | – |
| ModaVerse | I+V+A+TI+V+A+T | ImageBind | Linear Projector | LLaMA-2 | MLP | I: Stable Diffusion; V: Videofusion; A: AudioLDM | – | – |
| MM-Interleaved | I+TI+T | I: CLIP ViT-L/14 | Cross-attention | Vicuna-13B | Tiny Transformer | I: Stable Diffusion-2.1 | – | – |
@@ 3.2 MM IT
MM IT is a method that entails fine-tuning of pre-trained MM-LLMs using instruction-formatted datasets Wei et al. (2021). Through this process, MM-LLMs can generalize to unseen tasks by adhering to new instructions, thereby enhancing zero-shot performance. This straightforward yet impactful concept has catalyzed subsequent success in the field of NLP, exemplified by works such as InstructGPT Ouyang et al. (2022), OPT-IML Iyer et al. (2022), and InstructBLIP Dai et al. (2023).
@@
MM IT comprises Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), aiming to align with human intents and enhance the interaction capabilities of MM-LLMs. SFT converts part of the PT stage data into an instruction-aware format. Using visual Question-Answer (QA) as an example, various templates may be employed like (1) "<Image>{Question}" A short answer to the question is; (2) "<Image>" Examine the image and respond to the following question with a brief answer: "{Question}. Answer:"; and so on. Next, it fine-tunes pre-trained MM-LLMs using the same optimization objectives. SFT datasets can be structured as either single-turn QA or multi-turn dialogues.
@@
After SFT, RLHF involves further fine-tuning of the model, relying on feedback regarding the MM-LLMs’ responses (e.g., Natural Language Feedback (NLF) labeled manually or automatically) Sun et al. (2023b). This process employs a reinforcement learning algorithm to effectively integrate the non-differentiable NLF. The model is trained to generate corresponding responses conditioned on the NLF Chen et al. (2023i); Akyürek et al. (2023). The statistics for SFT and RLHF datasets are presented in Table 4 of Appendix G.
@@ 4 SOTA MM-LLMs
As shown in Figure 3, we classify the SOTA MM-LLMs from both functional and design perspectives. In the design division, “Tool-using” denotes treating the LLM as black box and providing access to certain MM expert systems to perform specific MM tasks via reasoning, while “End-to-End” signifies that the entire model is trained jointly in an end-to-end manner. Based on the previously defined design formulations, we also conduct a comprehensive comparison of the architectures and training dataset scales for of these SOTA MM-LLMs, as illustrated in Table 1. Next, we will summarize their developmental trends and briefly introduce the core contributions of some representative models in Appendix E.
@@ 5 Benckmarks and Performance
To offer a comprehensive performance comparison, we have compiled a table featuring major MM-LLMs across Vision-Language (VL) benchmarks gathered from various papers Li et al. (2023e); Chen et al. (2023d, f); Lin et al. (2023), shown in Table 2. The information of these benchmarks can be found in Appendix F. Next, we will extract training recipes that boost the effectiveness of MM-LLMs, drawing insights from SOTA models.
@@
Firstly, higher image resolution can incorporate more visual details for the model, benefiting tasks that require fine-grained details. For example, LLaVA-1.5 and VILA employ a resolution of , while Qwen-VL and MiniGPT-v2 utilize . However, higher resolutions lead to longer token sequences, incurring additional training and inference costs. MiniGPT-v2 addresses this by concatenating adjacent visual tokens in the embedding space to reduce length. Recently, Monkey Li et al. (2023l) proposed a solution to enhance the resolution of input images without retraining a high-resolution visual encoder, utilizing only a low-resolution visual encoder, supporting resolutions up to . To enhance the understanding of rich-text images, tables, and document content, DocPedia Feng et al. (2023) introduced a method to increase the visual encoder resolution to , overcoming the limitations of poorly performing low resolutions in open-sourced ViT. Secondly, the incorporation of high-quality SFT data can significantly improve performance in specific tasks, as evidenced by the addition of ShareGPT4V data to LLaVA-1.5 and VILA-13B, as shown in Table 2. Moreover, VILA reveals several key findings: (1) Performing PEFT on the LLM Backbone promotes deep embedding alignment, crucial for ICL; (2) Interleaved Image-Text data proves beneficial, whereas Image-Text pairs alone are sub-optimal; (3) Re-blending text-only instruction data (e.g., unnatural instruction Honovich et al. (2022)) with image-text data during SFT not only addresses the degradation of text-only tasks but also enhances VL task accuracy.